OpenStudy (anonymous):

How do i find the coordinates of x + 3y = 9 ?

OpenStudy (anonymous):

coordinates of what?

OpenStudy (anonymous):

For a graph

OpenStudy (anonymous):

solve for y it will give you slope intercept form then plot

OpenStudy (anonymous):

Whats x?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

ok, let me explain. x+3y = 9 // subtract x from both sides to remove the x from the left side of the equation. 3y = 9-x // divide both sides by 3 y = 3-1/3(x) // move the x left on the right side of the equal sign. y = (-1/3)x+3